sand drops vertically at rate 2kg/sec on a conveyer belt moving with 0.2m/s horizontally. what is extra force required and rate of change of kinetic energy?

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Extra force required = 2*0.2=0.4 N....

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sangeeta
5 years ago
We know that, F=dp/dt
And dp=d(mv)=vdm
Therefore, F=vdm/dt (dm/dt= 2kg/sec..Given)
F= 0.2×2= 0.4
•°• F=0.4 Newton